7) Suppose the demand for pizza in a small isolated town is p = 10 - Q. There are only two firms, A and B, and each has a cost function TC = 2 + q. Determine the Cournot equilibrium.
Answer #1

P = 10-Q P = 10- (@, +(22) [Lince there are two froms] P = 10-a, az TR, = 100, -Q, ² - 6,22 TR2 = 102-Q, Qq - Q² MR = 10-20,